OpenStudy (anonymous):

If Pr[C] = 0.52, Pr[D] = 0.52, and Pr[CuD] = 0.82. What is Pr[C|D] and Pr[D|C] ?

OpenStudy (amistre64):

well, we should have enough to calculate Pr[CnD] for the given parts

OpenStudy (amistre64):

do you recall any formulas that may help us?

OpenStudy (amistre64):

recall that P(A|B) is asking for the amount of A that is a part of B  so its just the amount associated with AnB comapred to B

OpenStudy (anonymous):

I just don't know how to get Pr[CnD] and that's what I need to solve the equation @am

OpenStudy (amistre64):

well, would you agree that P(AuB) would amount to the whole venn diagram? if so, then if we do P(A) + P(B) we have actually doubled some values ... the AnB part  A = (1+2), B=(2+3) A+B = 1+2+2+3 ---- we need to subtract the overlap (AnB)